• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name aniol

    Log entry time 19:18:07 on June16,2004

    Entry number 125985

    keyword=septa status

    The right septum curent is still ramping up but it is near its proper value. The left septum helium level is at 4.23%. It does not seem to have changed for the last 3 hours. Todd Ewing (tech on call ) says we should wait longer since there is helium going into the reservoir. When the lef helium level gets to the set value we can click the button to "automatic". The left reservoir is now on "manual". There is very little beam now because of the Moller measurement and there is no target. The left septum is still below 7.75K, however, we should watch its temperature once we put beam on the helium target. We can not turn on the left septum until the helium reservoir is at the proper level.